# EmoFace: Emotion-Content Disentangled Speech-Driven 3D Talking Face with Mesh Attention

### v1

@@ Methods
As mentioned before, in order to fuse the prediction results of the two branches, we propose a novel temporal-spatial attention module Mesh Attention, as shown in Figure 2. Mesh Attention analyses the emotion features and content features to get the weight of both features. Firstly, $F_{e1}\in\mathbb{R}^{B\times T\times C}$ and $F_{c1}\in\mathbb{R}^{B\times T\times C}$ go through one layer of MLP to get $F_{e1}\in\mathbb{R}^{B\times T\times C/4}$ and $F_{c1}\in\mathbb{R}^{B\times T\times C/4}$ with lower channel dimensions, and then they are concatenated in channel dimensions, followed by another layer of MLP to further reduce the channel number to get the fusion feature $F_{audio}\in\mathbb{R}^{B\times T\times C/16}$. Secondly, we perform the vertex position embedding of the FLAME topology to obtain the position embedding feature $F_{vertices-emb}\in\mathbb{R}^{B\times V\times C/16}$. Furthermore, to obtain the temporal features of the mesh vertices, we integrate the fusion feature $F_{audio}$ and the position embedding feature $F_{vertices-emb}$. Specifically, the fusion feature $F_{audio}$ is unsqueezed in the second dimension and then repeated $V$ times to obtain $F^{\prime}_{audio}\in\mathbb{R}^{B\times V\times T\times C/16}$. Similarly, $F_{vertices-emb}$ is unsqueezed in the first dimension and then repeated $T$ times to obtain $F^{\prime}_{vertices-emb}\in\mathbb{R}^{B\times V\times T\times C/16}$. The obtained features are summed eventually to get the integrated temporal features of the vertices $F\in\mathbb{R}^{B\times V\times T\times C/16}$.
@@
In purpose of further exploring the temporal-spatial connection between the vertices, we design a new graph-based convolution operator, the 3D spiral convolution (SpiralConv3D). In spatial domain, it determines the convolution centre and generates a series of enumerated vertices based on adjacency, followed by 1-ring vertices, 2-ring vertices, and so on, until all vertices containing k rings are included. SpiralConv3D determines the adjacency as follows:
$0{-ring}(v)$ $={v}$
$(k+1){-}ring(v)$ $=\mathcal{N}(k{-ring}(v))\setminus k{-disk}(v)$
$k{-disk}(v)$ $=\cup_{i=0,...,k}i{-ring}(v)$ (2)
@@
$connect_{\delta}(V_{ti})=\left\{\begin{aligned} \{V_{t-\delta+1,i},\ldots,V_{t-1,i},V_{t,i})\}&,t\geq\delta\\ \{Pad(\delta-t),V_{1,i},\ldots,V_{t,i}\}&,t<\delta\end{aligned}\right.$ (3)
@@
where $V_{ti}$ is the $i$-th vertex of moment $t$, $0\leq t\leq T$ and $0\leq i\leq 5023$. Meanwhile, $connect_{\delta}(V_{ti})$ denotes the connection between $V_{ti}$ and the vertices of past $\delta$ frames, and padding is used when $t<\delta$. Similar to 3D ConvNets (Tran et al. 2015), SpiralConv3D considers both temporal and spatial correlations. In contrast, SpiralConv3D applies this spatiotemporal modeling to mesh sequences with a sparse spatial distribution and a dense temporal distribution, considering convolution as the use of fully connected layers for feature fusion:
$SpiralConv3D(v)=W(f(connect_{\delta}(k{-disk}(v))))+b$ (4)
@@
where $f$ denotes the feature extractor, $W$ and $b$ are learnable weights and bias. To our knowledge this work exploits 3D GraphConv in the context of supervised training datasets and modern deep architectures to achieve the best performance on 3D facial animation.
Training and Testing
Figure 4: Qualitative comparison of the facial movements of the different methods on 3D-RAVDESS(left) and VOCASET(right). On RAVDESS, we generate facial animations of saying the sentence “Kids are talking by the door.” with surprised. On VOCA-Test, we generate facial animations of saying the sentence “How many crystal modifications of uranium hydride are extent?” without emotion. EmoFace generates more realistic facial movements that match the speech, whether it’s emotional or not.
During the training phase, we adopt a novel self-growing scheme instead of teacher-forcing or autoregression scheme. This scheme means using teacher-forcing scheme in the first stage and scheduled sampling scheme in the second stage. Self-growing scheme only fully guides in the initial epochs and then gradually reduces the guidance after the artificially selected epoch to enhance the robustness of the model. In our experiments, we observe that the training process of the self-growing scheme is smoother, allowing the model to gradually accept its own outputs, which weakens the impact of error accumulation and enhances the generation quality. We follow the self-reconstruction loss, cross-reconstruction loss, velocity loss and classification loss designed by EmoTalk (Peng et al. 2023b) as follows:
${L}_{self}=\left\|M_{c1,e1}-\hat{M}_{c1,e1}\right\|,{L}_{cross}=\left\|M_{c1,e2}-\hat{M}_{c1,e2}\right\|,$ (5)
${L}_{vel}=\left\|M^{t}-M^{t-1},\hat{M}^{t}-\hat{M}^{t-1}\right\|,$ (6)
${L}_{cls}=-\sum_{i}\sum_{c=1}^{N_{e}}\left(y_{ic}*\log p_{ic}\right),$ (7)
where $N_{e}$ represents the number of distinct emotion categories, $y_{ic}$ is the observation function that determines whether sample $i$ carries the emotion label $c$, and $p_{ic}$ denotes the predicted probability that sample $i$ belongs to class $c$. Especially, we include intermediate supervision for the predictions of the two independent branches as well, which we call branch supervision:
${L}_{branch}$ $=\left\|M^{e}_{c1,e1},\hat{M}_{c1,e1}\right\|+\left\|M^{c}_{c1,e1},\hat{M}_{c1,e1}\right\|$
$+\left\|M^{e}_{c1,e2},\hat{M}_{c1,e2}\right\|+\left\|M^{c}_{c1,e2},\hat{M}_{c1,e2}\right\|.$ (8)
Since in the second stage we use the schedule sampling scheme, we will get a second group of losses, denoted as $L^{\prime}_{self}$, $L^{\prime}_{cross}$, $L^{\prime}_{vel}$ and $L^{\prime}_{branch}$. The overall function is given by:
$L$ $=L_{1}+L_{2}+\lambda_{5}L_{class}$
$=\lambda_{1}(L_{self}+L^{\prime}_{self})+\lambda_{2}(L_{cross}+L^{\prime}_{cross})$
$+\lambda_{3}(L_{vel}+L^{\prime}_{vel})+\lambda_{4}(L_{branch}+L^{\prime}_{branch})+\lambda_{5}L_{class}$
@@
where $\lambda_{1}=\lambda_{2}=\lambda_{4}=1000$, $\lambda_{3}=500$ and $\lambda_{5}=0.0001$ in all of our experiments. During inference, EmoFace autoregressively predicts the mesh sequence corresponding to previous 3D talking faces. That is, at each moment $t$, EmoFace predicts the face motion $M_{t}$ conditioned on the raw audio $A$, the prior sequence of face motions $M_{\leq t}$ , the speaking style $s_{n}$, and the emotion level $l_{n}$. The $s_{n}$ and $l_{n}$ are determined by the speaker, and thus altering the one-hot identity and intensity vectors can manipulate the output in different styles and emotional levels.
Experiments
Experimental Settings
@@
We employed two datasets, including the non-emotional dataset VOCASET and the emotional dataset 3D-RAVDESS, where the 3D-RAVDESS dataset is the reconstructed dataset that we constructed.
@@
VOCASET dataset. VOCASET consists of 480 face mesh sequences from 12 subjects. Each mesh sequence is 60fps and is between 3 and 4 seconds long in duration. Meanwhile, each 3D face mesh has 5023 vertices. We follow the data configuration of VOCA for a fair comparison.
@@
3D-RAVDESS Dataset. The RAVDESS (Livingstone and Russo 2018) is a multimodal emotion recognition dataset. The dataset contains 24 actors (12 male and 12 female), each actor has 60 sentences with a total of 1440 face mesh sequences and corresponding speech. Each actor in this dataset provides data on different emotion categories and emotion levels, including neutral, calm, happy, sad, angry, fearful, disgusted, and surprised. In this case, the data corresponding to the first 20 subjects of the dataset are used for training, the 21st and 22nd subjects are used for validation, and the last two subjects are used for testing. We first process 1440 videos from the original RAVDESS dataset to convert the frame rate to 30 fps. Then, the emotional 3D faces are reconstructed using EMOCA (Daněček, Black, and Bolkart 2022) to obtain a sequence of 5023 mesh vertices under the FLAME network topology. Due to the jittery results, we use Kalman Filter on the FLAME parameters and fix the last three of the pose parameters to obtain 3D head mesh sequences with smooth front-face view. Our 3D-RAVDESS dataset consists of these mesh sequences and the corresponding speech from the original 2D dataset.
Training details. Our method takes speech data as input and also provides mesh templates, emotion levels, and speaker IDs as conditions. The sampling rate of the speech is $16kHz$ and the frame rate of the mesh sequence is 30 frames per second. During training, the model is end-to-end optimised using the Adam optimizer. The learning rate and batch size are set to $10^{-4}$ and $1$ respectively. The model is trained on a single NVIDIA V100 with 200 epochs on VOCASET and 225 epochs on 3D-RAVDESS.
Baseline Methods. Our algorithm is compared with FaceFormer, CodeTalker, EmoTalk and Selftalk and achieve state-of-the-art performance on both VOCASET and 3D-RAVDESS datasets. Since the official EmoTalk is a blendshape-based method that is incompatible with the marked facial mesh vertices provided by the datasets, we modify its output layer to directly predict the offsets of the vertices. Our method, FaceFormer and CodeTalker require conditions on a training speaker identity during inference. Therefore, for unseen subjects in the test dataset, we obtain the predictions of them by conditioning on all training identities. To adapt the input of non-silent frame to the 3D-RAVDESS dataset, we use the first frame as a condition in the training of EmoFace, FaceFormer and CodeTalker. All of these models are autoregressive, facilitating the use of input data from the first frame, while SelfTalk and EmoTalk are not autoregressive models.
@@
FaceFormer 4.1090 0.4858 5.8462 1.4149
CodeTalker 3.9445 0.5074 15.6160 3.4675
EmoTalk 3.7798 0.4862 6.6076 1.5994
SelfTalk 3.2238 0.4562 6.2560 1.5685
EmoFace(Ours) 2.8669 0.4664 5.0343 1.0196
Table 1: Quantitative evaluation results on 3D-RAVDESS and VOCASET.(For better visualization, we scale up the LVE and EVE by a factor of $10^{-5}$.)
Quantitative evaluation
To evaluate the lip synchronisation, we compute the lip vertex error (LVE) that is used in previous work (Fan et al. 2022). This metric computes the maximum $\ell_{2}$ error among all lip vertices in the test set and averages $\ell_{2}$ error across all frames. Additionally, the emotional vertex error (EVE) (Peng et al. 2023b) is used to reflect the full emotional expression. It measures the maximum $\ell_{2}$ error of all eye and forehead vertices in the test set and averages $\ell_{2}$ error of them. Table 1 shows that our algorithm achieves lower LVE and EVE on both datasets. In particular, our LVE and EVE on 3D-RAVDESS dataset is 20% and 35% lower than SelfTalk, respectively, demonstrating significant advantages of our algorithm in handling emotional speech-driven 3D face.
Qualitative evaluation
We qualitatively compare the driving effects between different models. The left side of Figure 4 shows the driving results of different models for the same speech with strong surprised emotion, and the right side shows the driving results of different models in non-emotional speech. In terms of lip synchronisation, we use red boxes in Figure 4 to mark the locations where the different algorithms show significant differences. EmoFace shows a greater amplitude of movement than the other algorithms, as can be seen especially in the shape of the lips for “are”, “talking” and “door”, which provides a better reflection of surprise and is more consistent with real movements of the lips. In addition, it also has a higher degree of mouth closure than other methods, for example, when pronouncing “many”, “crystal” and “extent”. Furthermore, the green box in Figure 4 indicates that EmoFace drives more obvious and natural expressions. To further demonstrate the superiority of our algorithm for both emotional and non-emotional audio inputs, a supplementary video is provided for more detailed comparisons. Incidentally, since we also encountered that CodeTalker crashes when trained on RAVDESS, as mentioned in DF-3DFace (Park et al. 2023) and LaughTalk (Sung-Bin et al. 2024), we have to use the results under the weights trained on VOCASET for zero-shot prediction.
Ablation experiment
We make a comparison of ablation experiments on the 3D-RAVDESS dataset. We train our method using teacher-forcing withtout self-growing and get higher LVE and EVE, as shown in Table 2. We believe the reason is that guiding is too strong, leading to poor robustness, and the prediction errors in the previous frames accumulate and affect the prediction of subsequent frames. However, with gradually weakened guiding in the self-growing scheme, EmoFace is trained to take into account the errors of the previous frames, which is similar to the situation during inference.
{tblr}
row3 = c, cell22 = c, cell23 = c, cell42 = c, cell43 = c, cell31 = l, hline1,5 = -0.08em, hline2 = -0.05em, Modules & LVE (×$\times 10^{-5}$mm) EVE (×$\times 10^{-5}$mm)

### v3

@@ 3.2 Mesh Attention
Firstly, $F_{e1}\in\mathbb{R}^{B\times T\times C}$ and $F_{c1}\in\mathbb{R}^{B\times T\times C}$ go through one layer of MLP to get $F_{e1}\in\mathbb{R}^{B\times T\times C/4}$ and $F_{c1}\in\mathbb{R}^{B\times T\times C/4}$ with lower channel dimensions, and then they are concatenated in channel dimensions, followed by another layer of MLP to further reduce the channel number to get the fusion feature $F_{audio}\in\mathbb{R}^{B\times T\times C/16}$. Secondly, we perform the vertex position embedding of the FLAME topology to obtain the position embedding feature $F_{vertices-emb}\in\mathbb{R}^{B\times V\times C/16}$. Furthermore, to obtain the temporal features of the mesh vertices, we integrate the fusion feature $F_{audio}$ and the position embedding feature $F_{vertices-emb}$. Specifically, the fusion feature $F_{audio}$ is unsqueezed in the second dimension and then repeated $V$ times to obtain $F^{\prime}_{audio}\in\mathbb{R}^{B\times V\times T\times C/16}$. Similarly, $F_{vertices-emb}$ is unsqueezed in the first dimension and then repeated $T$ times to obtain $F^{\prime}_{vertices-emb}\in\mathbb{R}^{B\times V\times T\times C/16}$. The obtained features are summed eventually to get the integrated temporal features of the vertices $F\in\mathbb{R}^{B\times V\times T\times C/16}$.
Figure 2: Self-growing scheme. In the first stage, the model inputs all groundtruth frames at once and directly predicts the next frame corresponding to each input frame. In the second stage, the input is changed to a fusion of the groundtruth frames and the predicted frames from the previous stage, and the same prediction process is applied to obtain the final prediction results.
@@
In purpose of further exploring the temporal-spatial connection between vertices, we design a new graph-based convolution operator, 3D spiral convolution (SpiralConv3D). In spatial domain, it determines the convolution centre and generates a series of enumerated vertices based on adjacency, followed by 1-ring vertices, 2-ring vertices, and so on, until all vertices containing k rings are included. SpiralConv3D determines the adjacency as follows:
$0{-ring}(v)$ $={v},$
$(k+1){-}ring(v)$ $=\mathcal{N}(k{-ring}(v))\setminus k{-disk}(v),$
$k{-disk}(v)$ $=\cup_{i=0,...,k}i{-ring}(v),$ (1)
@@
$cnt_{\delta}(V_{ti})=\left\{\begin{aligned} \{V_{t-\delta+1,i},\ldots,V_{t-1,i},V_{t,i})\}&,t\geq\delta\\ \{Pad(\delta-t),V_{1,i},\ldots,V_{t,i}\}&,t<\delta\end{aligned}\right.,$ (2)
@@
where $V_{ti}$ is the $i$-th vertex of moment $t$, $0\leq t\leq T$ and $0\leq i\leq 5023$. Meanwhile, $cnt_{\delta}(V_{ti})$ denotes the connection between $V_{ti}$ and the vertices of past $\delta$ frames, and padding is used when $t<\delta$. Similar to 3D ConvNets Tran et al. (2015), SpiralConv3D considers both temporal and spatial correlations. In contrast, SpiralConv3D applies this spatio-temporal modeling to mesh sequences with a sparse spatial distribution and a dense temporal distribution, considering convolution as the use of fully connected layers for feature fusion:
$S\!piralConv3D(v)=$ $W(f(cnt_{\delta}(k{-disk}(v))))$
$+b,$ (3)
@@
where $f$ denotes the feature extractor, $W$ and $b$ are learnable weights and bias. To our knowledge, this work exploits 3D GraphConv in the context of supervised training datasets and modern deep architectures to achieve the best performance on 3D facial animation.
3.3 Training and Testing
3.3.1 Self-growing Scheme
During the training phase, we adopt a novel self-growing scheme instead of teacher-forcing or autoregression scheme, as shown in Figure 2. Self-growing scheme divides the training process in each epoch into two stages. In the first stage, the model inputs all groundtruth frames at once, and then uses Temporal Bias Fan et al. (2022) to eliminate the influence of future frames and directly predicts the next frame corresponding to each input frame. In the second stage, the model adopts the fusion strategy to obtain new input data and repeats the prediction operation in the previous stage. The fusion strategy is adapted according to the training process. Specifically, in the first few $\theta$ epochs, we keep all the groundtruth frames $GT$ as inputs; in the remaining epochs, the first half gradually replace the groundtruth frames $GT$ with the predicted frames $P$ from the previous stage using the cosine function, and the second half directly abandon the groundtruth frames. The fusion strategy can be denoted as follows:
$I_{n,t}=\left\{\begin{array}[]{ll}GT_{t},n<\theta\text{\enspace}or\text{\enspace}rand()<cos(\frac{\pi(n-\theta)}{N-\theta})\\ P_{t},n>\frac{(N+\theta)}{2}\text{\enspace}or\text{\enspace}rand()\geq cos(\frac{\pi(n-\theta)}{N-\theta})\end{array},\right.$ (6)
where $n$ and $t$ denote the $n$-th epoch and the $t$-th frame, respectively. $N$ means the total number of epochs and $rand()$ generates a random number from a uniform distribution over the interval [0, 1]. Self-growing scheme fully guides in the initial epochs and then gradually reduces the guidance and accepts its own outputs, which weakens the impact of error accumulation and improves the robustness and generation quality.
@@ 3.3.1 Self-growing Scheme
During inference, EmoFace autoregressively predicts the mesh sequence corresponding to previous 3D talking faces. Specifically, at each moment $t$, EmoFace predicts the face motion $M_{t}$ conditioned on the raw audio $A$, the prior sequence of face motions $M_{\leq t}$ , the speaking style $s_{n}$, and the emotion level $l_{n}$ at each moment. The $s_{n}$ and $l_{n}$ are determined by the speaker, and thus altering the one-hot identity and intensity vectors can manipulate the output in different styles and emotional levels.
3.3.2 Training Strategy of Emotion-Content Disentanglement Module
We set a pseudo-training pair of samples with the same content but different emotions. The emotion features and content features are extracted respectively and the two emotion features are exchanged as input of transformer decoder to implement cross reconstruction. Since that we expect both the emotion branch and the content branch to be as strong as possible in modeling the speech-mesh mapping relationship, our emotion branch and content branch make separate predictions for the mesh offsets of the next frame. We incorporate the intermediate supervision into the predictions of two separate branches. However, both branches have significant limitations in their respective abilities to model facial motions. On the one hand, although the features extracted from the content branch are strongly correlated with the lips, the inability to extract long-term emotional features and the lack of the type and intensity of emotion are not conducive to predicting the magnitude of mouth opening. On the other hand, though the features extracted from the emotion branch are strongly correlated with the motions of the eyes and the face expressions, the lack of short-term content features results in insufficiently accurate predictions of the lips. Thus, as mentioned before, Mesh Attention is designed to fuse the driving results of the two branches in temporal and spatial domains to obtain the final prediction. Our self-reconstruction loss, cross-reconstruction loss, velocity loss and classification loss are as follows:
${L}_{self}$ $=\left\|M^{e}_{c1,e1}-\hat{M}_{c1,e1}\right\|_{1}+\left\|M^{c}_{c1,e1}-\hat{M}_{c1,e1}\right\|_{1}$
$+\left\|M_{c1,e1}-\hat{M}_{c1,e1}\right\|_{1},$ (7)
${L}_{cross}$ $=\left\|M^{e}_{c1,e2}-\hat{M}_{c1,e2}\right\|_{1}+\left\|M^{c}_{c1,e2}-\hat{M}_{c1,e2}\right\|_{1}$
$+\left\|M_{c1,e2}-\hat{M}_{c1,e2}\right\|_{1},$ (8)
${L}_{vel}$ $=\left\|M^{t}_{c1,e1}-M^{t-1}_{c1,e1},\hat{M}^{t}_{c1,e1}-\hat{M}^{t-1}_{c1,e1}\right\|_{1}$
$+\left\|M^{t}_{c1,e2}-M^{t-1}_{c1,e2},\hat{M}^{t}_{c1,e2}-\hat{M}^{t-1}_{c1,e2}\right\|_{1},$ (9)
${L}_{cls}$ $=-\sum_{i}\sum_{\phi=1}^{N_{e}}\left(y_{i\phi}*\log p_{i\phi}\right),$ (10)
where $M_{cx,ey}$ represents predicted mesh sequence with content $x$ and emotion $y$. $M^{c}$ and $M^{e}$ represents predicted mesh sequence of content branch and emotion branch, respectively. $M^{t}$ means the $t$-th frame of mesh sequence, $\hat{M}$ means the groundtruth, $N_{e}$ represents the number of distinct emotion categories, $y_{i\phi}$ is the observation function that determines whether sample $i$ carries the emotion label $\phi$, and $p_{i\phi}$ denotes the predicted probability that sample $i$ belongs to class $\phi$. Note that the first two terms of self-reconstruction loss and cross-reconstruction loss are intermediate supervision of the two branches. The overall function is given by:
$L=\lambda_{1}L_{self}+\lambda_{2}L_{cross}+\lambda_{3}L_{vel}+\lambda_{4}L_{cls},$ (11)
where $\lambda_{1}=\lambda_{2}=1000$, $\lambda_{3}=500$ and $\lambda_{4}=0.0001$ in all of our experiments.
4 Experiments
4.1 Experimental Settings
@@
We employ two datasets, including the non-emotional dataset VOCASET and the emotional dataset 3D-RAVDESS, where the 3D-RAVDESS dataset is the dataset we constructed.
@@
VOCASET dataset. VOCASET Cudeiro et al. (2019) consists of 480 face mesh sequences from 12 subjects. Each mesh sequence is 60fps and is between 3 and 4 seconds long in duration. Meanwhile, each 3D face mesh has 5023 vertices. We follow the data configuration of VOCA for a fair comparison.
@@
3D-RAVDESS Dataset. The RAVDESS Livingstone and Russo (2018) is a multimodal emotion recognition dataset. As shown in Table 1, the dataset contains 24 actors (12 male and 12 female), each actor has 60 sentences with a total of 1440 face mesh sequences and corresponding speech. Each actor in this dataset provides data on different emotion categories and emotion levels, including neutral, calm, happy, sad, angry, fearful, disgusted, and surprised. In this case, the data corresponding to the first 20 subjects of the dataset are used for training, the 21st and 22nd subjects are used for validation, and the last two subjects are used for testing. We first process 1440 videos from the original RAVDESS dataset to convert the frame rate to 30 fps. Then, the emotional 3D faces are reconstructed using EMOCA Daněček et al. (2022) to obtain a sequence of 5023 mesh vertices under the FLAME network topology. Due to the jittery results, we use Kalman Filter Kalman (1960) on the FLAME parameters and fix the last three of the pose parameters to obtain 3D head mesh sequences with smooth front view. Our 3D-RAVDESS dataset consists of these mesh sequences and the corresponding speech from the original 2D dataset.
Subject Gender Ratio Text Emotion Sentence FPS Topology Training:Validation:Testing
24 1:1 60 8 1440 30 FLAME 10:1:1
Table 1: Statistics of 3D-RAVDESS dataset.
Figure 3: Qualitative comparison of the facial movements of the different methods on 3D-RAVDESS (left) and VOCASET (right). On 3D-RAVDESS, we generate facial animations of saying the sentence “Kids are talking by the door.” with surprised. On VOCA-Test, facial animations of saying the sentence “How many crystal modifications of uranium hydride are extent?” without emotion are generated. Significant differences in the lip region are denoted by red boxes. EmoFace generates more realistic facial movements that match the speech, whether it’s emotional or not.
4.2 Quantitative evaluation
To evaluate the lip synchronisation, we compute the lip vertex error (LVE) that is used in previous work Fan et al. (2022). This metric computes the maximum $\ell_{2}$ error among all lip vertices in the test set and averages $\ell_{2}$ error across all frames. Additionally, the emotional vertex error (EVE) Peng et al. (2023b) is used to reflect the full emotional expression. It measures the maximum $\ell_{2}$ error of all eye and forehead vertices in the test set and averages $\ell_{2}$ error of them. Table 2 demonstrates significant advantages of our algorithm in handling emotional speech-driven 3D face. In particular, our LVE and EVE on 3D-RAVDESS dataset is 20% and 35% lower than SelfTalk, respectively.
@@ 4.2 Quantitative evaluation
VOCA Cudeiro et al. (2019) 4.9245
MeshTalk Richard et al. (2021) 4.5441
FaceFormer Fan et al. (2022) 4.1090 0.4858 5.8462 1.4149
CodeTalker Xing et al. (2023) 3.9445 0.5074 15.6160 3.4675
EmoTalk Peng et al. (2023b) 3.7798 0.4862 6.6076 1.5994
SelfTalk Peng et al. (2023a) 3.2238 0.4562 6.2560 1.5685
TalkingStyle Song et al. (2024) 3.5245 0.4686 5.9737 1.4471
EmoFace(Ours) 2.8669 0.4664 4.8863 0.9509
Table 2: Quantitative evaluation results on VOCASET and 3D-RAVDESS.(For better visualization, we scale up the LVE and EVE by a factor of $10^{-5}$.)
4.3 Qualitative evaluation
We qualitatively compare the driving effects of cross-identity between different models by driving the new 3D character templates in the test set with the speaking style of the training characters. The left side of Figure 3 shows the driving results of different models for the same speech with strong surprised emotion, and the right side shows the driving results of different models in non-emotional speech. Since that the speaking style of the new characters is unknown in the test data, the model may be driven slightly differently to the real results. In terms of lip synchronisation, EmoFace shows a greater amplitude of movement, which is particularly evident in the lip shapes for “are", “talking" and “door". This provides a more accurate reflection of surprise and is more consistent with real movements of the lips. In addition, it also has a higher degree of mouth closure than other methods, for example, when pronouncing “many", “crystal" and “extent". Furthermore, the full face comparison in Figure 3 indicates that EmoFace drives more obvious and natural expressions. To further demonstrate the superiority of our algorithm for both emotional and non-emotional audio inputs, a supplementary video is provided for more detailed comparisons.
4.4 Ablation experiments
In this section, we conduct ablation experiments to investigate the impact of our training strategy and Mesh Attention. All ablation experiments are conducted on the 3D-RAVDESS dataset.
Modules
LVE
(×$\times 10^{-5}$mm)
EVE
(×$\times 10^{-5}$mm)